# TSI-Bench: Benchmarking Time Series Imputation

### v1

@@ 5 Conclusion
In this paper, we introduce TSI-Bench, a comprehensive benchmark for time series imputation. Our evaluation of 28 algorithms on 8 real-world datasets demonstrates that successful data imputation depends on various factors, including the deep learning structures, missingness rates and patterns across space-time, and the datasets themselves. Although this process is complex, our benchmark suite and experimental results can serve as a reference for future research, and provide practical imputation guidelines in real scenarios.
We envision TSI-Bench as a long-term evolving project and are committed to its continuous development. Our future roadmap includes integrating more advanced deep-learning models and expanding the range of datasets. Additionally, we aim to enhance the user-friendliness of TSI-Bench, with the goal of establishing it as a prevalent tool for time series imputation.
Acknowledgments and Disclosure of Funding
@@ A.1 Datasets for evaluation
BeijingAir ItalyAir PeMS Pedestrian ETTh1 Electricity PhysioNet2012 PhysioNet2019
@@ A.1.1 Air quality datasets
  • BeijingAir [^1] https://archive.ics.uci.edu/dataset/501/beijing+multi+site+air+quality+data [51]: BeijingAir dataset includes hourly air pollutants data from 12 nationally-controlled air-quality monitoring sites from March 2013 to February 2017. The dataset contains 420,768 instances, with 6 main air pollutants and 6 relevant meteorological variables at multiple sites in Beijing. We aggregate these variables from all sites together so this dataset has 12*12=132 features.
  • ItalyAir [^2] https://archive.ics.uci.edu/dataset/360/air+quality [52]: ItalyAir dataset contains 9358 instances of hourly averaged responses of 5 metal oxides from chemical sensors, along with hourly averaged concentrations of 7 pollutants from certified analyzers, from March 2004 to February 2005.
@@ A.1.2 Traffic datasets
  • PeMS [^3] https://PeMS.dot.ca.gov: The PeMS is a collection of hourly data from California Department of Transportation, which describes the road occupancy rates measured by different sensors on San Francisco Bay area freeways.
  • Pedestrian [^4] https://www.timeseriesclassification.com/description.php?Dataset=MelbournePedestrian [53]: The City of Melbourne, Australia has developed an automated pedestrian counting system to better understand pedestrian activity within the municipality, The data of a specific region is from the whole year of 2017.
@@ A.1.4 Healthcare datasets
  • PhysioNet2012777https://physionet.org/content/challenge-2012/1.0.0/ [54]: The PhysioNet/Computing in Cardiology Challenge 2012 dataset includes 12,000 ICU patient records from the MIMIC II Clinical database, version 2.6 [70], focusing on patient-specific prediction of in-hospital mortality using data from the first 48 hours of ICU admission. Not all variables are available in all cases, hence about 80% values are missing in this dataset. The whole dataset has three subsets, and we only use the subset A in our experiments.
  • PhysioNet2019888https://physionet.org/content/challenge-2019/1.0.0/ [55]: The PhysioNet Challenge 2019 dataset includes clinical data from ICU patients across three hospitals, with a total of 40,336 patient records and 40 clinical and physiological variables for each patient. Note that this dataset has two subsets, and we only use the subset A in our setting.
@@ A.2 Datasets Preprocessing Details
BeijingAir, ItalyAir, PeMS, Electricity, and ETT_h1 are all long time series continuously collected from certain sources. Therefore, to generate them into the train, validation, and test sets and to avoid data leakage, we should first split them according to the time period. In BeijingAir, the first 28 months (2013/03 - 2015/06) of data are taken for training, the following 10 months (2015/07 - 2016/04) are for validation, and the left 10 months (2016/05 - 2017/02) are for test. In PeMS, the training set tasks the first 15 months (2016/07 - 2017/09) of data, and the validation set and test set take the following 4 months (2017/10 - 2018/01) and 6 months (2018/02 - 2018/07) respectively. Electricity uses the first 10 months (2011/01 - 2011/10) as the test set, the following 10 months (2011/11 - 2012/08) for validation, and the last 28 months (2012/09 - 2014/12) for training. The training, validation, and test sets of ETT_h1 separately take the first 14 months (2016/07 - 2017/08), the following 5 months (2017/09 - 2018/01), and the last 5 months (2018/02 - 2018/06). ItalyAir is split into 60%, 20%, and 20% for training, validation, and test. The sliding window function is applied to these five datasets to produce data samples. The window size of ItalyAir is 12, and 24 for BeijingAir and PeMS, 48 for ETT_h1, 96 for Electricity. The sliding length is kept the same as the window size to guarantee there is no overlap between generated samples.
@@ Appendix B Detailed Description of Models in TSI-Bench
Table 4: Summary of the 28 benchmarked algorithms, separated by Architecture
@@
iTransformer [60] Forecasting Transformer ICLR 2024
SAITS [20] Imputation Transformer ESWA 2023
Nonstationary [59] Forecasting Transformer NeurIPS 2022
ETSformer [58] Forecasting Transformer ArXiv 2022
PatchTST [33] Forecasting Transformer ICLR 2023
Crossformer [43] Forecasting Transformer ICLR 2022
Informer [42] Forecasting Transformer AAAI 2021
Autoformer [31] Forecasting Transformer NeurIPS 2021
Pyraformer [57] Forecasting Transformer ICLR 2021
Transformer [56] Forecasting Transformer NeurIPS 2017
BRITS [19] Imputation RNN NeurIPS 2018
MRNN [61] Imputation RNN TBME 2018
GRUD [62] Imputation RNN Scientific reports 2018
TimesNet [13] Generic CNN ICLR 2023
MICN [63] Forecasting CNN ICLR 2022
SCINet [64] Forecasting CNN NeurIPS 2022
StemGNN [34] Forecasting GNN NeurIPS 2020
FreTS [67] Forecasting MLP NeurIPS 2024
Koopa [66] Forecasting MLP NeurIPS 2024
DLinear [44] Forecasting MLP AAAI 2023
FiLM [65] Forecasting MLP NeurIPS 2022
CSDI [30] Imputation Diffusion NeurIPS 2021
US-GAN [68] Imputation GAN AAAI 2021
GP-VAE [28] Imputation VAE AISTATS 2020
@@ B.1 Transformer-based models
  • iTransformer [60]: iTransformer repurposes the Transformer architecture by applying attention and feed-forward networks to inverted dimensions, embedding time points into variate tokens to better capture multivariate correlations and nonlinear representations, achieving state-of-the-art performance in time series forecasting.
  • SAITS [20]: SAITS (Self-Attention-based Imputation for Time Series) is a self-attention imputation transformer with a weighted combination of two diagonally-masked self-attention (DMSA) blocks. It is designed to handle missing data in time series, ensuring robust and accurate data imputation.
  • Nonstationary [59]: Nonstationary (short for Nonstationary Transformer) addresses the challenge of non-stationarity in time series sequences by incorporating adaptive components that adjust to varying statistical properties over time.
  • ETSformer [58]: ETSformer integrates exponential smoothing methods with Transformer models, aiming to provide accurate time series forecasting by combining statistical and deep learning approaches.
  • PatchTST [33]: PatchTST uses a patching strategy combined with a Transformer architecture to enhance the time series forecasting task by capturing both local and global patterns effectively.
  • Crossformer [43]: Crossformer leverages cross-dimensional attention to model intricate dependencies within multivariate time series data, achieving good performance in complex forecasting application scenarios.
  • Informer [42]: Informer enhances efficiency in long time series forecasting by employing a self-attention distillation mechanism, which reduces redundant information while maintaining forecasting accuracy.
  • Autoformer [31]: Autoformer introduces a novel decomposition architecture with an auto-correlation mechanism, effectively capturing both seasonal and trend patterns in time series forecasting tasks.
  • Pyraformer [57]: Pyraformer is designed for long-term time series forecasting, utilizing a pyramid attention structure that efficiently captures temporal dependencies at multiple scales.
  • Transformer [71]: Transformer introduces the self-attention mechanism, which enables the processing of sequential data by attending to different positions within the sequence simultaneously, leading to significant advancements in natural language processing and time series fields.
@@ B.2 RNN-based models
  • BRITS [19]: BRITS (Bidirectional Recurrent Imputation for Time Series) employs a bidirectional recurrent imputation strategy to handle missing values in time series, improving forecasting accuracy through iterative refinement.
  • MRNN [61]: The Multi-directional Recurrent Neural Network (MRNN) is designed for estimating missing values in spatiotemporal data and time series by leveraging temporal dependencies and multi-directional information sequence.
  • GRUD [62]: GRUD (Gated Recurrent Unit for Decay) enhances the Gated Recurrent Unit (GRU) by incorporating decay mechanisms that model the impact of missing values over time, offering robust time series analysis.
@@ B.3 CNN-based models
  • TimesNet [13]: TimesNet is designed to efficiently model temporal patterns in time series data by incorporating multi-scale temporal convolutions and attention mechanisms. It can be used for short- and long-term forecasting, imputation, classification, and anomaly detection tasks.
  • MICN [63]: MICN (Multi-scale Inception Convolutional Network) is a convolutional network architecture that captures multi-scale temporal features and combines local and global context for more accurate time series forecasting.
  • SCINet [64]: SCINet introduces a novel recursive downsample-convolve-interact architecture for time series forecasting, leveraging multiple convolutional filters to extract and aggregate valuable temporal features from downsampled sub-sequences, improving forecasting accuracy over existing models.
@@ B.4 GNN-based models
  • StemGNN [34]: StemGNN (Spectral Temporal Graph Neural Network) introduces an approach for time series forecasting by jointly capturing inter-series correlations and temporal dependencies in the spectral domain using Graph Fourier Transform (GFT) and Discrete Fourier Transform (DFT), eliminating the need for pre-defined priors.
@@ B.5 MLP-based models
  • FreTS [67]: FreTS uses MLPs in the frequency domain for time series forecasting, leveraging global view and energy compaction properties to enhance forecasting performance by transforming time-domain signals and learning frequency components.
  • Koopa [66]: Koopa introduces a novel Koopman forecaster that disentangles time-variant and time-invariant components using Fourier Filter and employs Koopman operators for linear dynamics portrayal, achieving end-to-end forecasting with significant improvements in training time and memory efficiency.
  • DLinear [44]: DLinear introduces LTSF-Linear, a set of simple one-layer linear models, which outperform complex Transformer-based models in long-term time series forecasting by effectively preserving temporal information that Transformers inherently lose due to their permutation-invariant self-attention mechanism.
  • FiLM [65]: FiLM (Frequency improved Legendre Memory) introduces a novel approach by using Legendre Polynomials for historical information approximation, Fourier projection for noise reduction, and a low-rank approximation for computational efficiency, resulting in significant improvements in long-term forecasting accuracy.
@@ B.6 Generative models
  • CSDI [30]: CSDI (Conditional Score-based Diffusion Imputation) leverages conditional score-based diffusion models for accurate imputation and generation of missing values in time series applications.
  • US-GAN [68]: US-GAN integrates a classifier in a semi-supervised generative adversarial network to enhance its imputation of missing values in multivariate time series, leveraging both observed data and label information. Also, it introduces a temporal matrix to improve the discriminator’s ability to differentiate between observed and imputed components.
  • GP-VAE [28]: GP-VAE (Gaussian Process Variational Autoencoder) proposes a novel deep sequential latent variable model that combines VAE with a structured variational approximation to achieve non-linear dimensionality reduction and imputation, providing interpretable uncertainty estimates and improved imputation smoothness.
@@ Appendix C Details in TSI-Bench
C.1 Missing Patterns

### v2

@@ 5 Conclusion
In this paper, we introduce TSI-Bench, a comprehensive benchmark for time series imputation. Our evaluation encompasses 28 algorithms across 8 real-world datasets, illustrating that successful imputation hinges on factors like deep learning architectures, missing rates, missing patterns across space-time, and the datasets themselves. Additionally, incorporating uncertainty quantification in the evaluation process is essential, as it enables an assessment of confidence in the imputed data and helps to identify the limitations and reliability of the results. However, uncertainty quantification in time series imputation remains an emerging area, limited by the complexity of sequential dependencies, high-dimensional features, and the lack of standardized metrics.
Despite the complexities, TSI-Bench provides valuable insights and practical guidelines for time series imputation in real-world scenarios. We envision TSI-Bench as a continually evolving project. Our future plans include the integration of more advanced deep learning models, the deployment of more equitable evaluation metrics, and the expansion of our dataset collection. Additionally, we are committed to improving the user-friendliness of TSI-Bench, aiming to establish it as the standard tool for time series imputation.
6 Acknowledgments
@@ A.1.1 Air Quality Datasets
  • BeijingAir [^1] https://archive.ics.uci.edu/dataset/501/beijing+multi+site+air+quality+data (Zhang et al., 2017): BeijingAir dataset includes hourly air pollutants data from 12 nationally-controlled air-quality monitoring sites from March 2013 to February 2017. The dataset contains 11 continuous variables at multiple sites in Beijing. We aggregate these variables from all sites together so this dataset has 12*11=132 features.
  • ItalyAir [^2] https://archive.ics.uci.edu/dataset/360/air+quality (De Vito et al., 2008): ItalyAir dataset contains 9358 instances of hourly averaged responses of 5 metal oxides from chemical sensors, along with hourly averaged concentrations of 7 pollutants from certified analyzers, from March 2004 to February 2005.
@@ A.1.2 Traffic Datasets
  • PeMS [^3] https://PeMS.dot.ca.gov: The PeMS is a collection of hourly data from the California Department of Transportation, which describes the road occupancy rates measured by different sensors on San Francisco Bay area freeways.
  • Pedestrian [^4] https://www.timeseriesclassification.com/description.php?Dataset=MelbournePedestrian (Tang et al., 2020): The City of Melbourne, Australia has developed an automated pedestrian counting system to better understand pedestrian activity within the municipality, The data of a specific region is from the whole year of 2017.
@@ A.1.4 Healthcare Datasets
  • PhysioNet2012777https://physionet.org/content/challenge-2012/1.0.0/ (Silva et al., 2012): The PhysioNet/Computing in Cardiology Challenge 2012 dataset includes 12,000 ICU patient records from the MIMIC II Clinical database, version 2.6 (Saeed et al., 2011), focusing on patient-specific prediction of in-hospital mortality using data from the first 48 hours of ICU admission. Not all variables are available in all cases, hence about 80% values are missing in this dataset. The whole dataset has three subsets, and we only use the subset A in our experiments.
  • PhysioNet2019888https://physionet.org/content/challenge-2019/1.0.0/ (Reyna et al., 2020): The PhysioNet Challenge 2019 dataset includes clinical data from ICU patients across three hospitals, with a total of 40,336 patient records and 40 clinical and physiological variables for each patient. Note that this dataset has two subsets, and we only use the subset A in our setting.
@@ A.2 Datasets Preprocessing Details
BeijingAir, ItalyAir, PeMS, Electricity, and ETT_h1 are all long-time series continuously collected from certain sources. Therefore, to generate them into the train, validation, and test sets and to avoid data leakage, we should first split them according to the time period. In BeijingAir, the first 28 months (2013/03 - 2015/06) of data are taken for training, the following 10 months (2015/07 - 2016/04) are for validation, and the left 10 months (2016/05 - 2017/02) are for test. In PeMS, the training set tasks the first 15 months (2016/07 - 2017/09) of data, and the validation set and test set take the following 4 months (2017/10 - 2018/01) and 6 months (2018/02 - 2018/07) respectively. Electricity uses the first 10 months (2011/01 - 2011/10) as the test set, the following 10 months (2011/11 - 2012/08) for validation, and the last 28 months (2012/09 - 2014/12) for training. The training, validation, and test sets of ETT_h1 separately take the first 14 months (2016/07 - 2017/08), the following 5 months (2017/09 - 2018/01), and the last 5 months (2018/02 - 2018/06). ItalyAir is split into 60%, 20%, and 20% for training, validation, and test. The sliding window function is applied to these five datasets to produce data samples. The window size of ItalyAir is 12, and 24 for BeijingAir and PeMS, 48 for ETT_h1, and 96 for Electricity. The sliding length is kept the same as the window size to guarantee there is no overlap between generated samples.
@@ Appendix B Detailed Description of Models in TSI-Bench
Table 5: Summary of the 28 benchmarked algorithms, separated by Architecture.
@@
iTransformer (Liu et al., 2023b) Forecasting Transformer ICLR 2024
SAITS (Du et al., 2023) Imputation Transformer ESWA 2023
Nonstationary (Liu et al., 2022c) Forecasting Transformer NeurIPS 2022
ETSformer (Woo et al., 2022) Forecasting Transformer ArXiv 2022
PatchTST (Nie et al., 2023) Forecasting Transformer ICLR 2023
Crossformer (Zhang & Yan, 2022) Forecasting Transformer ICLR 2022
Informer (Zhou et al., 2021) Forecasting Transformer AAAI 2021
Autoformer (Wu et al., 2021) Forecasting Transformer NeurIPS 2021
Pyraformer (Liu et al., 2021) Forecasting Transformer ICLR 2021
Transformer (Yıldız et al., 2022) Forecasting Transformer NeurIPS 2017
BRITS (Cao et al., 2018) Imputation RNN NeurIPS 2018
MRNN (Yoon et al., 2018) Imputation RNN TBME 2018
GRUD (Che et al., 2018b) Imputation RNN Scientific reports 2018
TimesNet (Wu et al., 2023) Generic CNN ICLR 2023
MICN (Wang et al., 2022) Forecasting CNN ICLR 2022
SCINet (Liu et al., 2022a) Forecasting CNN NeurIPS 2022
StemGNN (Cao et al., 2020) Forecasting GNN NeurIPS 2020
FreTS (Yi et al., 2024) Forecasting MLP NeurIPS 2024
Koopa (Liu et al., 2024) Forecasting MLP NeurIPS 2024
DLinear (Zeng et al., 2023) Forecasting MLP AAAI 2023
FiLM (Zhou et al., 2022) Forecasting MLP NeurIPS 2022
CSDI (Tashiro et al., 2021) Imputation Diffusion NeurIPS 2021
US-GAN (Miao et al., 2021a) Imputation GAN AAAI 2021
GP-VAE (Fortuin et al., 2020) Imputation VAE AISTATS 2020
@@ B.1 Transformer-based Models
  • iTransformer (Liu et al., 2023b): iTransformer repurposes the Transformer architecture by applying attention and feed-forward networks to inverted dimensions, embedding time points into variate tokens to better capture multivariate correlations and nonlinear representations, achieving state-of-the-art performance in time series forecasting.
  • SAITS (Du et al., 2023): SAITS (Self-Attention-based Imputation for Time Series) is a self-attention imputation transformer with a weighted combination of two diagonally-masked self-attention (DMSA) blocks. It is designed to handle missing data in time series, ensuring robust and accurate data imputation.
  • Nonstationary (Liu et al., 2022c): Nonstationary (short for Nonstationary Transformer) addresses the challenge of non-stationarity in time series sequences by incorporating adaptive components that adjust to varying statistical properties over time.
  • ETSformer (Woo et al., 2022): ETSformer integrates exponential smoothing methods with Transformer models, aiming to provide accurate time series forecasting by combining statistical and deep learning approaches.
  • PatchTST (Nie et al., 2023): PatchTST uses a patching strategy combined with a Transformer architecture to enhance the time series forecasting task by capturing both local and global patterns effectively.
  • Crossformer (Zhang & Yan, 2022): Crossformer leverages cross-dimensional attention to model intricate dependencies within multivariate time series data, achieving good performance in complex forecasting application scenarios.
  • Informer (Zhou et al., 2021): Informer enhances efficiency in long-time series forecasting by employing a self-attention distillation mechanism, which reduces redundant information while maintaining forecasting accuracy.
  • Autoformer (Wu et al., 2021): Autoformer introduces a novel decomposition architecture with an auto-correlation mechanism, effectively capturing both seasonal and trend patterns in time series forecasting tasks.
  • Pyraformer (Liu et al., 2021): Pyraformer is designed for long-term time series forecasting, utilizing a pyramid attention structure that efficiently captures temporal dependencies at multiple scales.
  • Transformer (Vaswani et al., 2017): Transformer introduces the self-attention mechanism, which enables the processing of sequential data by attending to different positions within the sequence simultaneously, leading to significant advancements in natural language processing and time series fields.
@@ B.2 RNN-based Models
  • BRITS (Cao et al., 2018): BRITS (Bidirectional Recurrent Imputation for Time Series) employs a bidirectional recurrent imputation strategy to handle missing values in time series, improving forecasting accuracy through iterative refinement.
  • MRNN (Yoon et al., 2018): The Multi-directional Recurrent Neural Network (MRNN) is designed for estimating missing values in spatiotemporal data and time series by leveraging temporal dependencies and multi-directional information sequence.
  • GRUD (Che et al., 2018b): GRUD (Gated Recurrent Unit for Decay) enhances the Gated Recurrent Unit (GRU) by incorporating decay mechanisms that model the impact of missing values over time, offering robust time series analysis.
@@ B.3 CNN-based Models
  • TimesNet (Wu et al., 2023): TimesNet is designed to efficiently model temporal patterns in time series data by incorporating multi-scale temporal convolutions and attention mechanisms. It can be used for short- and long-term forecasting, imputation, classification, and anomaly detection tasks.
  • MICN (Wang et al., 2022): MICN (Multi-scale Inception Convolutional Network) is a convolutional network architecture that captures multi-scale temporal features and combines local and global context for more accurate time series forecasting.
  • SCINet (Liu et al., 2022a): SCINet introduces a novel recursive downsample-convolve-interact architecture for time series forecasting, leveraging multiple convolutional filters to extract and aggregate valuable temporal features from downsampled sub-sequences, improving forecasting accuracy over existing models.
@@ B.4 GNN-based Models
  • StemGNN (Cao et al., 2020): StemGNN (Spectral Temporal Graph Neural Network) introduces an approach for time series forecasting by jointly capturing inter-series correlations and temporal dependencies in the spectral domain using Graph Fourier Transform (GFT) and Discrete Fourier Transform (DFT), eliminating the need for pre-defined priors.
@@ B.5 MLP-based Models
  • FreTS (Yi et al., 2024): FreTS uses MLPs in the frequency domain for time series forecasting, leveraging global view and energy compaction properties to enhance forecasting performance by transforming time-domain signals and learning frequency components.
  • Koopa (Liu et al., 2024): Koopa introduces a novel Koopman forecaster that disentangles time-variant and time-invariant components using Fourier Filter and employs Koopman operators for linear dynamics portrayal, achieving end-to-end forecasting with significant improvements in training time and memory efficiency.
  • DLinear (Zeng et al., 2023): DLinear introduces LTSF-Linear, a set of simple one-layer linear models, which outperform complex Transformer-based models in long-term time series forecasting by effectively preserving temporal information that Transformers inherently lose due to their permutation-invariant self-attention mechanism.
  • FiLM (Zhou et al., 2022): FiLM (Frequency improved Legendre Memory) introduces a novel approach by using Legendre Polynomials for historical information approximation, Fourier projection for noise reduction, and a low-rank approximation for computational efficiency, resulting in significant improvements in long-term forecasting accuracy.
@@ B.6 Generative Models
  • CSDI (Tashiro et al., 2021): CSDI (Conditional Score-based Diffusion Imputation) leverages conditional score-based diffusion models for accurate imputation and generation of missing values in time series applications.
  • US-GAN (Miao et al., 2021a): US-GAN integrates a classifier in a semi-supervised generative adversarial network to enhance its imputation of missing values in multivariate time series, leveraging both observed data and label information. Also, it introduces a temporal matrix to improve the discriminator’s ability to differentiate between observed and imputed components.
  • GP-VAE (Fortuin et al., 2020): GP-VAE (Gaussian Process Variational Autoencoder) proposes a novel deep sequential latent variable model that combines VAE with a structured variational approximation to achieve non-linear dimensionality reduction and imputation, providing interpretable uncertainty estimates and improved imputation smoothness.
@@ Appendix C Details in TSI-Bench
C.1 Missingness
The concept of missingness can be understood through two fundamental aspects: missing mechanisms and missing patterns. Missing mechanisms describe the underlying processes that lead to missing values, while missing patterns define how these missing values manifest within the dataset.
Missing Mechanisms
The seminal taxonomy introduced by (Rubin, 1976) categorizes missing data mechanisms into three distinct types: Missing Completely At Random (MCAR), Missing At Random (MAR), and Missing Not At Random (MNAR). These classifications hinge on the conditional probability distribution $p(\mathbf{M}|\mathbf{\tilde{X}},\bm{\phi})$, where $\bm{\phi}$ parameterizes the stochastic mapping from the dataset $\mathbf{\tilde{X}}$ to the mask indicator $\mathbf{M}$. Specifically:
  • MCAR implies that the probability of data being missing is independent of observations and missingness. This can be expressed as: $p(\mathbf{M}|\mathbf{\tilde{X}},\bm{\phi})=p(\mathbf{M}|\bm{\phi})$.
  • MAR suggests that the missingness depends solely on observations, i.e., $p(\mathbf{M}|\mathbf{\tilde{X}},\bm{\phi})=p(\mathbf{M}|\mathbf{X},\bm{\phi})$.
  • MNAR indicates that the missingness is related to the unobserved data and potentially the observed data as well, i.e., $p(\mathbf{M}|\mathbf{\tilde{X}},\bm{\phi})=p(\mathbf{M}|\mathbf{\tilde{X}},\bm{\phi})$.
Missing Patterns
Despite the clarity the above taxonomy provides, it often falls short in capturing the complexity of missing data scenarios in modern datasets (Mitra et al., 2023). This limitation has driven the development of new frameworks (Mitra et al., 2023; Cini et al., 2022; Khayati et al., 2020) that better align with diverse real-world data.
To effectively analyze missingness in time series data, we employ three distinct patterns of missing data: point, subsequence, and block. These patterns are specifically chosen to represent typical missing data scenarios in time series analysis, providing a robust framework for evaluating different imputation strategies. The missingness construction functions are directly available in the Python library PyGrinder [^9] https://github.com/WenjieDu/PyGrinder. The visualization of them is plotted in figures LABEL:fig:point_missing_pattern, LABEL:fig:subseq_missing_pattern, and LABEL:fig:block_missing_pattern for vivid illustration.